# Few-Shot Scenario Testing for Autonomous Vehicles Based on Neighborhood Coverage and Similarity

### v1

@@ Abstract
Testing and evaluating the safety performance of autonomous vehicles (AVs) is essential before the large-scale deployment. Practically, the acceptable cost of testing specific AV model can be restricted within an extremely small limit because of testing cost or time. With existing testing methods, the limitations imposed by strictly restricted testing numbers often result in significant uncertainties or challenges in quantifying testing results. In this paper, we formulate this problem for the first time the “few-shot testing” (FST) problem and propose a systematic FST framework to address this challenge. To alleviate the considerable uncertainty inherent in a small testing scenario set and optimize scenario utilization, we frame the FST problem as an optimization problem and search for a small scenario set based on neighborhood coverage and similarity. By leveraging the prior information on surrogate models (SMs), we dynamically adjust the testing scenario set and the contribution of each scenario to the testing result under the guidance of better generalization ability on AVs. With certain hypotheses on SMs, a theoretical upper bound of testing error is established to verify the sufficiency of testing accuracy within given limited number of tests. The experiments of the cut-in scenario using FST method demonstrate a notable reduction in testing error and variance compared to conventional testing methods, especially for situations with a strict limitation on the number of scenarios.
@@
Few-shot testing, autonomous vehicles, scenario coverage, testing scenario library
@@ I Introduction
The fast development and experimental application of high-level autonomous vehicles (AVs) on open road underscores the crucial need for testing and evaluating safety performance to facilitate the large-scale deployment. However, with seemingly endless traffic scenarios in real world[1] and the low efficiency of testing rare events (e.g. accidents)[2], the available testing cost is far from meeting the practical requirements. Consequently, how to generate reliable testing results within the confines of a restricted testing number limit becomes a significant problem.
@@
In practical terms, the acceptable cost of testing specific AV model can be restricted within an extremely small limit. For third-party testing organizations and governmental bodies, the generation of an extensive array of testing scenarios for all potential AV models, particularly during open road testing, is not pragmatic. Besides, with the rapid iterative development of autonomous driving technique, conducting a thorough evaluation of AV performance within the research and development cycle becomes increasingly infeasible. Consequently, there exists a compelling need to generate quick and accurate testing results of AV using the smallest possible testing numbers or within a strictly limited set of scenarios. Moreover, the quantitative and explainable results are needed as a foundational benchmark for comparing the performance of diverse AVs, which create additional difficulties for testing with limited costs. Remarkably, we term this problem the “few-shot testing” (FST) problem in this paper, marking the first instance of defining and addressing this specific issue to the best of our knowledge.
@@
Although a lot of efforts have been made to search for a smaller testing scenario library or accelerate the testing process, the problem of FST remains unsolved. As a practical method, some autonomous driving companies keep a scenario library from logged data or expert knowledge to verify the reliability of their AVs before on-road deployment[3]. Searching for critical scenarios is a commonly used scheme to generate a smaller testing scenario library[4]. Based on knowledge[5], scenario clustering[6, 7], scenario coverage[8], optimization strategy[9, 10] or some well-designed models[11, 12, 13, 14], many methods are capable of generating a representative scenario library with certain risks. However, the efficiency of these methods is usually measured by the proportion of critical scenarios or some pre-defined reasonable criteria. The quantitative result for the performance index of AV is hard to be acquired.
@@
Statistical methods represent an effective approach for quantifying the performance index of AV model while generating critical scenarios to accelerate the testing process[15, 16, 17, 18]. Based on naturalistic driving data (NDD) and naturalistic driving environment (NDE), the performance of AVs can be estimated with a critical distribution. However, although unbiased testing results can be acquired through sampling, many critical scenarios may be similar or reduplicative. Additionally, controlling the testing variance in a smaller test set becomes challenging due to uncertainty, thereby diminishing the effectiveness of statistical methods in few-shot testing (FST) scenarios.
@@
In this paper, we propose systematically the FST framework to tackle the problem of quantifying the performance of AV with smallest possible or strictly limited testing numbers. Meanwhile, A testing error bound with regard to the restricted number of testing scenarios can be provided to verify the accuracy of FST method. In order to maintain the theoretical advantages and quantify the performance indeces of AV, we generate testing scenarios based on NDD and NDE. Furthermore, to mitigate the large uncertainty of NDE-based statistical methods associated with the restricted testing numbers, we frame the FST problem as an optimization problem and iteratively search for FST scenario set for an accurate and reliable testing result. In the context of fixed few-shot scenarios and an unknown AV model before testing, the essence of FST method lies not in the direct selection of critical scenarios but in identifying a set of scenarios with optimal generalization ability based on SMs. This is similar to the generalization ability concept employed in classic few-shot learning methods [19, 20].
@@
Following the basic idea of FST framework, we introduce a dynamic neighborhood coverage set and a similarity measurement for scenarios in the test set. According to the limitation on testing numbers and specific scenario samples, the confidence and coverage of scenarios are dynamically adjusted in the optimization process under the guidance of the generalization ability or the estimation error bound. With the gradient descent method, a fixed small set of scenarios can be selected with a minimized upper bound of error. The performance index of real AV model could be quantified with definite accuracy with some hypothesss on SMs. In order to validate the FST method, we conducted a case study in the cut-in scenario. Compared with commonly used testing methods, the average error and variance of FST is significantly smaller. Notably, we find that the accuracy of FST method is less impaired with a smaller size of testing set compared with the other methods.
@@ II-A Performance Index Quantification with NDE
where $\mathcal{X}$ is the set of all possible scenarios. With statistics given by NDD, the probability measure of specific scenario can be derived as $p(x),x\in\mathcal{X}$, which is the exposure frequency of the scenario in real world. Define $X$ as the corresponding random variable and then we have $X\sim p(x)$. Provided that $A$ is the event of interest during testing (usually accidents for AVs), the performance index of AVs can be calculated as
$\mu=P(A)$ $=E_{p}[P(A|X)]$ (2)
$=\sum_{x\in\mathcal{X}}{P(A|x)\cdot p(x)},$
where $\mu=P(A)$ is the performance index with respect to event $A$ in NDE and $P(A|x)$ is the performance measure of the vehicle under test in the scenario $x$. For the purpose of quantifying AV performance and ensuring the theoretical testing accuracy, NDE is adopted in this paper.
@@ II-B Statistical Testing Problem
$\tilde{\mu}_{CMC}=\frac{1}{n}\sum_{i=1}^{n}{P(A|x_{i})},X_{i}\sim p(x),$ (3)
where $\mathcal{X}_{s}$ is generated from natural distribution $p(x)$ of NDE and $n$ could increase dynamically throughout the testing process. When $A$ denotes accidents, the accident rate of AV is generally a tiny value, so CMC faces the problem of “curse of rarity” and exhibits extremely low efficiency and accuracy[2]. Consequently, it is practically impossible to evaluate AVs with CMC coupled with a small number of tests.
Towards addressing this issue, importance sampling (IS) is proposed to accelerate the testing process[17, 18, 15, 16], the strategy for IS to quantify accident rate is
$\tilde{\mu}_{IS}=\frac{1}{n}\sum_{i=1}^{n}{P(A|x_{i})\cdot\frac{p(x_{i})}{q(x_{i})}},X_{i}\sim q(x),$ (4)
where $\mathcal{X}_{n}$ is generated from an important distribution $q(x)$. The testing efficiency of IS can be much higher than CMC. However, in practice the estimation result of IS could not converge to an accurate value for a strictly limited number of tests.
@@ II-C Few-Shot Testing Problem
In this paper, we focus on the situation when a limited size of scenario set $\mathcal{X}_{s}$ is given as $n$, denoting $\mathcal{X}_{s,n}\triangleq\{x_{i}\in\mathcal{X},i=1,2,...,n\}$. As $n$ is a small number (e.g. $n=5,10$), our primary emphasis is on the estimation error caused by specific testing set $\mathcal{X}_{s,n}$ and the goal is to minimize the evaluation error on specific AV model, that is
$\mathop{\min}_{\mathcal{X}_{s,n}}Err=|\tilde{\mu}-\mu|,$ (5)
@@
$\mathop{\min}_{\mathcal{X}_{s,n}}Err=|f(P(A|x_{1}),...,P(A|x_{n}))-\mu|.$ (7)
Statistical testing method like CMC and IS have good theoretical properties that the estimation is unbiased. However, as scenarios are sampled from probabilistic distribution, the uncertainty is hard to be controlled with few testing numbers and the estimation errors is possible to be a large value. In Eq. (7), all relevant notations are fixed value but not random variables, so the uncertainty of sampling is eliminated.
Towards solving the problem in Eq. (7), there are still several challenges remain:
(1).The estimation function $f$ is extremely flexible, which calls for a concrete and effective estimation strategy;
(2).As $P(A|x_{i})$ is obtained for specific AV model and $\mu$ is the ground truth of performance index for this AV, the optimal solution of Eq. (7) is relevant to the unknown AV model under test;
(3).If we have an approximation of $P(A|x_{i})$ or $\mu$ before testing, the unknown gap between real AV model and the approximation may directly affect the estimation error and result in a low accuracy.
@@ III-A Upper Bound of Estimation Error
AV plays an important role in the formulation (7). Because we do not know the information of AV model before testing, we use the surrogate models (SMs) to represent for the prior information of AV. Note that the generated FST scenario set $\mathcal{X}_{s,n}$ is fixed after the optimization is finished, but there are a diverse of possible AV model for test. Because the error between SMs and AVs under test may directly affect the estimation error, we suppose the information of SMs form a set $\mathcal{X}_{s,n}$. For each possible SM $m_{i}\in\mathcal{M}$, the performance measure $P_{i}(A|x)$ could be acquired through tests. Subsequently, under the hypothesis that real AV model satisfies $m^{*}\in\mathcal{M}$, the fixed error can be further written as an error bound
$Err\leq\mathop{\max}_{m_{i}\in\mathcal{M}}|\tilde{\mu}_{i}-\mu_{i}|,$ (8)
where $\tilde{\mu}_{i}$ and $\mu_{i}$ is respectively the accident rate and estimation of accident rate of $m_{i}$:
$\tilde{\mu}_{i}=f(P_{i}(A|x_{1}),...,P_{i}(A|x_{n})),\ \mu_{i}=E_{p}[P_{i}(A|X)].$ (9)
@@
$Err$ $\leq\mathop{\max}_{m_{i}\in\mathcal{M}}\{|\tilde{\mu}_{i}-\mu_{i}|\}+|\Delta\tilde{\mu}-\Delta{\mu}|$ (10)
@@ III-B Coverage and Similarity-Based Estimation Strategy
$f(P(A|x_{1}),...,P(A|x_{n}))=\sum_{i=1}^{n}{P(A|x_{i})\cdot w(x_{i};\mathcal{X}_{s,n})},$ (11)
where $w(x_{i};\mathcal{X}_{s,n})$ is the weight function of each scenario sample $x_{i}$ and we suppose that $\sum_{i=1}^{n}{w(x_{i};\mathcal{X}_{s,n})}=1$. Note from Eq. (11) that there is $w(x_{i};\mathcal{X}_{s,n})=1/n$ for CMC method and $w(x_{i};\mathcal{X}_{s,n})=p(x_{i})/[nq(x_{i})]$ for IS method. Traditional weight functions are determined only by certain sample $x_{i}$ and the size of test set $n$ hence lacking the global perspective. Besides, the $\mathcal{X}_{s,n}$ is sampled based on distributions. By making use of a relatively flexible form of weight functions and test set, the few-shot accuracy may be guaranteed.
Now we come to the discussion on the weight function. Following the idea of fully utilizing each valid testing scenario and avoiding redundant information, we want to take advantage of the concept of scenario coverage. The coverage of a scenario is hopeful to be a reliable measurement of its contribution to the testing result. Coverage is a commonly used and effective approach to verify the performance of AVs[22]. Experientially and intuitively, the similarity among scenarios allows us to use the performance of AV in a representative scenario as an approximation of its neighborhood.
Usually the coverage of scenario is deterministic with some pre-determined rules or models[23, 24]. As FST method tries to evaluate the performance of AV under fixed size of test set, the accuracy and expected contribution of each scenario sample for FST should change along with the size of test set. Therefore, we propose a dynamic neighborhood coverage set to adjust the relative coverage of scenarios selected for test dynamically with respect to $\mathcal{X}_{s,n}$. The coverage of one sample $x_{i}\in\mathcal{X}_{s,n}$ is decided by the similarity between all scenarios in the scenario space and the test set $\mathcal{X}_{s,n}$. We define the coverage set of $x_{i}$ by
$\mathcal{C}(x_{i};\mathcal{X}_{s,n})\triangleq\{x^{\prime}|\mathop{\arg\max}_{j=1,...,n}S(x^{\prime},x_{j})=i\}.$ (12)
$S(x^{\prime},x)$ is a similarity measurement which can be defined as the inverse of normalized Euclidean norm for simplicity $S(x^{\prime},x)\triangleq 1/||x-x^{\prime}||_{2}$. By adopting this brief form of similarity measurement, the boundary of coverage sets is actually a separate hyperplane in the scenario space. Then the weight of $x_{i}$ is computed with the sum of all scenarios in its coverage set weight by exposure frequency
@@
With the definition of coverage set and weight function, we can see that the estimation $\tilde{\mu}$ is continuous providing the continuity of scenario space $X,p(x)$ and performance measure function $P(A|x)$. With this property it is easy to prove that for specific AV model $m^{*}$ and $n\geq 2$, the optimal estimation error $\tilde{\mu}$ is 00 by selecting optimal $\mathcal{X}_{s,n}$. This ensures the theoretical optimality of FST method with exactly accurate prior knowledge. The real-world scenario variables are usually continuous but the discretization precision may affect the optimality to some extent.

### v2

@@ Abstract
Testing and evaluating the safety performance of autonomous vehicles (AVs) is essential before the large-scale deployment. Practically, the number of testing scenarios permissible for a specific AV is severely limited by tight constraints on testing budgets and time. With the restrictions imposed by strictly restricted numbers of tests, existing testing methods often lead to significant uncertainty or difficulty to quantifying evaluation results. In this paper, we formulate this problem for the first time the “few-shot testing” (FST) problem and propose a systematic framework to address this challenge. To alleviate the considerable uncertainty inherent in a small testing scenario set, we frame the FST problem as an optimization problem and search for the testing scenario set based on neighborhood coverage and similarity. Specifically, under the guidance of better generalization ability of the testing scenario set on AVs, we dynamically adjust this set and the contribution of each testing scenario to the evaluation result based on coverage, leveraging the prior information of surrogate models (SMs). With certain hypotheses on SMs, a theoretical upper bound of evaluation error is established to verify the sufficiency of evaluation accuracy within the given limited number of tests. The experiment results on cut-in scenarios demonstrate a notable reduction in evaluation error and variance of our method compared to conventional testing methods, especially for situations with a strict limit on the number of scenarios.
@@
Few-shot testing, autonomous vehicles, scenario coverage, testing scenario set
@@ I Introduction
The fast development and experimental application of high-level autonomous vehicles (AVs) on open road have brought about the crucial need and emerging research interests[1, 2, 3] on testing and evaluating their safety performance to facilitate the large-scale deployment. However, with seemingly endless traffic scenarios in real world[4] and the low efficiency of testing rare events (e.g. crashes)[5], the available testing budget is far from meeting the practical requirements. Consequently, how to generate reliable testing and evaluation results within the confines of a restricted limit of testing number becomes a significant problem.
@@
In practical terms, the acceptable budget of testing specific AV model can be restricted within an extremely small limit. For third-party testing organizations and governmental bodies, the generation of an extensive array of testing scenarios for all potential AV models, particularly during open road testing, is not pragmatic. Besides, with the rapid iterative development of autonomous driving technique, conducting a thorough evaluation of AV performance within the research and development cycle becomes increasingly infeasible. Consequently, there exists a compelling need to generate quickly the accurate evaluation results of AV using the smallest possible testing numbers or within a strictly limited set of scenarios. Moreover, the quantitative and explainable results are needed as a foundational benchmark for comparing the performance of diverse AVs, which create additional difficulties for testing with limited costs. Remarkably, we term this problem the “few-shot testing” (FST) problem in this paper, marking the first instance of defining and addressing this specific issue to the best of our knowledge.
@@
Although a lot of efforts have been made to search for a smaller testing scenario set or accelerated tests, the problem of FST remains unsolved. As a practical method, some autonomous driving companies keep a scenario set from logged data or expert knowledge to verify the reliability of their AVs before on-road deployment[6]. Searching for critical scenarios is a commonly used scheme to generate a smaller testing scenario set[7]. Based on knowledge[8], scenario clustering[9, 10], scenario coverage[11], optimization strategy[12, 13] or some well-designed models[14, 15, 16, 17], many methods are capable of generating a representative scenario set with certain risks. However, the efficiency of these methods is usually measured by the proportion of critical scenarios or some pre-defined reasonable criteria. The quantitative result for the performance index of AV is hard to be acquired.
@@
Statistical methods represent an effective approach for quantifying the performance index of AV model while generating critical scenarios to accelerate the testing process[18, 19, 20, 21, 22, 23, 24]. Based on naturalistic driving data (NDD) and naturalistic driving environment (NDE), the performance of AVs can be estimated with a critical distribution. However, although unbiased results can be acquired through sampling, many critical scenarios may be similar or reduplicative. Additionally, controlling the testing variance in a smaller test set becomes challenging due to uncertainty, thereby diminishing the effectiveness of these methods in FST scenarios.
@@
In this paper, we propose systematically the FST framework to tackle the problem of quantifying the performance of AV with smallest possible or strictly limited numbers of test. Meanwhile, A testing error bound with regard to the restricted number of testing scenarios can be derived to verify the accuracy of FST method. In order to maintain the theoretical advantages and quantify the performance indices of AV, we generate testing scenarios based on NDD and NDE. Furthermore, to mitigate the large uncertainty of NDE-based statistical methods associated with the restricted numbers of test, we frame the FST problem as an optimization problem and iteratively search for FST scenario set for an accurate and reliable testing and evaluation result. In the context of fixed few-shot scenarios and an unknown AV model before testing, the essence of FST method lies not in the direct selection of critical scenarios but in identifying a set of scenarios with optimal generalization ability based on SMs. This is similar to the generalization ability concept employed in classic few-shot learning methods [25, 26].
@@
Following the basic idea of FST framework, we introduce a dynamic neighborhood coverage set and a similarity measurement for scenarios in the test set. According to the limit on numbers of testing scenarios and specific scenario samples, the confidence and coverage of scenarios are dynamically adjusted in the optimization process under the guidance of the generalization ability or the estimation error bound. With the gradient descent method, a fixed small set of scenarios can be selected with a minimized upper bound of error. The performance index of real AV model could be quantified with definite accuracy with some hypothesis on SMs. In order to validate the FST method, we conducted a case study in the cut-in scenario. Compared with commonly used testing methods, the average error and variance of FST is significantly smaller. Notably, we find that the accuracy of FST method is less impaired with a smaller size of testing set compared with the other methods.
@@ II-A Performance Index Quantification with NDE
where $\mathcal{X}$ is the set of all possible scenarios. With statistics given by NDD, the probability measure of specific scenario can be derived as $p(x),x\in\mathcal{X}$, which is the exposure frequency of the scenario in real world. Define $X$ as the corresponding random variable and then we have $X\sim p(x)$. Provided that $A$ is the event of interest during testing (usually crashes for AVs), the performance index of AVs can be calculated as
$\mu=P(A)$ $=\mathbb{E}_{p}[P(A|X)]$ (2)
$=\sum_{x\in\mathcal{X}}{P(A|x)p(x)},$
where $\mu=P(A)$ is the performance index with respect to event $A$ in NDE and $P(A|x)$ is the performance measure of the vehicle under test in the scenario $x$.
@@ II-B Statistical Testing Problem
$\tilde{\mu}_{\rm{CMC}}=\frac{1}{n}\sum_{i=1}^{n}{P(A|x_{i})},X_{i}\sim p(x),$ (3)
where $\mathcal{X}_{s}$ is generated from natural distribution $p(x)$ of NDE and $n$ could increase dynamically throughout the testing procedure. When $A$ denotes crashes, the crash rate of AV is generally a tiny value and CMC faces the problem of “curse of rarity”, which results in extremely low efficiency and accuracy[5]. Consequently, it is practically impossible to evaluate AVs with CMC coupled with a small number of tests.
Towards addressing this issue, importance sampling (IS) is proposed to accelerate the testing process[21, 22, 18], the strategy for IS to quantify the crash rate is
$\tilde{\mu}_{\rm{IS}}=\frac{1}{n}\sum_{i=1}^{n}{P(A|x_{i})\frac{p(x_{i})}{q(x_{i})}},X_{i}\sim q(x),$ (4)
where $\mathcal{X}_{s}$ is generated from an important distribution $q(x)$. The testing efficiency of IS can be much higher than CMC under ideal condition. However, in practice the estimation result of IS could not converge to an accurate value for a strictly limited number of tests.
@@ II-C Few-Shot Testing Problem
In this paper, we focus on the situation where a limited size of scenario set $\mathcal{X}_{s}$ is given as $n$, denoting $\mathcal{X}_{s,n}\triangleq\{x_{i}\in\mathcal{X},i=1,2,...,n\}$. As $n$ is a small number (e.g. $n=5,10$), our primary emphasis is on the estimation error caused by specific testing set $\mathcal{X}_{s,n}$ and the goal is to minimize the evaluation error on specific AV model, that is
$\mathop{\min}_{\mathcal{X}_{s,n}}E=|\tilde{\mu}-\mu|,$ (5)
@@
$\mathop{\min}_{\mathcal{X}_{s,n}}E=|f(P(A|x_{1}),...,P(A|x_{n}))-\mu|.$ (7)
Statistical testing methods like CMC and IS have theoretical benefits that the estimation is unbiased. However, as scenarios are sampled from probabilistic distribution, the uncertainty is hard to be controlled with few testing numbers and the estimation errors is possible to be a large value. In Eq. (7), all relevant notations are fixed value but not random variables, so the uncertainty of sampling is eliminated.
Towards solving the problem in Eq. (7), there are still several challenges remaining:
(1) The estimation function $f$ is extremely flexible, which calls for a concrete and effective estimation strategy;
(2) As $P(A|x_{i})$ is obtained for specific AV model and $\mu$ is the ground truth of performance index for this AV, the optimal solution of Eq. (7) is relevant to the unknown AV model under test;
(3) If we have an approximation of $P(A|x_{i})$ or $\mu$ before testing, the unknown gap between real AV model and the approximation may directly affect the estimation error and result in a low accuracy.
@@ III-A Upper Bound of Estimation Error
AV plays an important role in the formulation (7). Because we do not know the information of AV model before testing, we use the surrogate models (SMs) to represent for the prior information of AV. Note that the generated FST scenario set $\mathcal{X}_{s,n}$ is fixed after the optimization is finished, but there are diverse possible AV models for test. Because the error between SMs and AVs under test may directly affect the estimation error, we suppose the information of SMs form a set $\mathcal{X}_{s,n}$. For each possible SM $m_{i}\in\mathcal{M}$, the performance measure $P_{i}(A|x)$ could be tested and acquired. Subsequently, under the hypothesis that real AV model satisfies $m^{*}\in\mathcal{M}$, the fixed error can be further written as an error bound
$E\leq\mathop{\max}_{m_{i}\in\mathcal{M}}|\tilde{\mu}_{i}-\mu_{i}|,$ (8)
where $\tilde{\mu}_{i}$ and $\mu_{i}$ is the crash rate and estimation of crash rate of $m_{i}$, respectively:
$\tilde{\mu}_{i}=f(P_{i}(A|x_{1}),...,P_{i}(A|x_{n})),\ \mu_{i}=\mathbb{E}_{p}[P_{i}(A|X)].$ (9)
@@
$E$ $\leq\mathop{\max}_{m_{i}\in\mathcal{M}}\{|\tilde{\mu}_{i}-\mu_{i}|\}+|\Delta\tilde{\mu}-\Delta{\mu}|$ (10)
@@ III-B Coverage and Similarity-Based Estimation Strategy
$f(P(A|x_{1}),...,P(A|x_{n}))=\sum_{i=1}^{n}{P(A|x_{i})w(x_{i};\mathcal{X}_{s,n})},$ (11)
where $w(x_{i};\mathcal{X}_{s,n})$ is the weight function of each scenario sample $x_{i}$ and we suppose that $\sum_{i=1}^{n}{w(x_{i};\mathcal{X}_{s,n})}=1$. Note from Eq. (11) that there is $w(x_{i};\mathcal{X}_{s,n})=1/n$ for CMC method and $w(x_{i};\mathcal{X}_{s,n})=p(x_{i})/[nq(x_{i})]$ for IS method. Traditional weight functions are determined only by certain sample $x_{i}$ and the size of test set $n$ hence lacking the global perspective. Besides, the $\mathcal{X}_{s,n}$ is sampled based on distributions, which brings about large uncertainty. By making use of a relatively flexible form of weight functions and test set, the few-shot accuracy may be guaranteed.
Now we come to the discussion on the weight function. Following the idea of fully utilizing each valid testing scenario and avoiding redundant information, we want to take advantage of the concept of scenario coverage. The coverage of a scenario is hopeful to be a reliable measurement of its contribution to the evaluation result. Coverage is a commonly used and effective approach to verify the performance of AVs[28]. Experientially and intuitively, the similarity among scenarios allows us to utilize the performance of AV in a representative scenario as an approximation of its neighborhood.
Usually the coverage of scenario is deterministic with some pre-determined rules or models[29, 30]. As FST method aims at evaluating the performance of AV under fixed size of test set, the accuracy and expected contribution of each scenario sample for FST should change along with the size of test set. Therefore, we propose a dynamic neighborhood coverage set to adjust the relative coverage of scenarios selected for test dynamically with respect to $\mathcal{X}_{s,n}$. The coverage of one sample $x_{i}\in\mathcal{X}_{s,n}$ is decided by the similarity between all scenarios in the scenario space and the test set $\mathcal{X}_{s,n}$. We define the coverage set of $x_{i}$ by
$\mathcal{C}(x_{i};\mathcal{X}_{s,n})\triangleq\left\{x^{\prime}|\mathop{\arg\max}_{j=1,...,n}S(x^{\prime},x_{j})=i\right\},$ (12)
where $S(x^{\prime},x)$ is a similarity measurement which can be defined as the inverse of normalized Euclidean norm for simplicity
$S(x^{\prime},x)\triangleq\frac{1}{||x-x^{\prime}||_{2}}.$ (13)
By adopting this brief form of similarity measurement, the boundary of coverage sets is actually a separate hyperplane in the scenario space. Then the weight of $x_{i}$ is computed with the sum of all scenarios in its coverage set weight by exposure frequency
@@
With the definition of coverage set and weight function, the estimation $\tilde{\mu}$ is continuous providing the continuity of scenario space $X,p(x)$ and performance measure function $P(A|x)$. With this property it is easy to prove that for specific AV model $m^{*}$ and $n\geq 2$, the optimal estimation error $\tilde{\mu}$ is 00 by selecting optimal $\mathcal{X}_{s,n}$. This ensures the theoretical optimality of FST method with exactly accurate prior knowledge. Specifically, the real-world scenario variables are usually continuous but the discretization precision may affect the optimality to some extent.